# Automatic Network Planning with Digital Radio Twin

### v2

@@ V-B Automatic Network Planning
The network performance, as defined in Eq.7, for all compared network planning methods is illustrated in Fig.7. As we can see, exhaustive search yields the highest overall performance, while random sampling performs significantly worse than both AutoPlan and ES. Notably, AutoPlan achieves performance nearly identical to that of ES, demonstrating its effectiveness in accurate and efficient BS deployment. In terms of computation time, ES requires $602$ minutes, approximately $58.33$ times longer than AutoPlan, highlighting the latter’s practical efficiency.
@@
Moreover, we evaluate the sensitivity of coverage to transmit power in Fig. 7, which plots the coverage achieved by AutoPlan as a function of the number of newly deployed BSs under different transmit power levels. As expected, coverage improves with increased power; however, the gain becomes marginal beyond $30,\mathrm{dBm}$. For example, when five new BSs are installed, increasing the power from $30,\mathrm{dBm}$ to $43,\mathrm{dBm}$ only results in an improvement of $2\%$. Considering the trade-off between performance and power consumption, we adopt $43,\mathrm{dBm}$ as the default setting in subsequent experiments. This value yields the highest observed coverage and corresponds to typical configurations of micro-base stations.
@@ VI Related Work
In contrast, AI/ML based methods aim to enhance scalability and adaptability by replacing manual engineering and handcrafted models with data-driven learning. AutoBS [17] employs reinforcement learning to select BS locations, enabling fast inference during deployment. However, it requires large-scale pretraining on labeled environment performance pairs, which is both computationally expensive and time-consuming to collect. OSSN [18], a Transformer-based framework, unifies radio map estimation and site selection to reduce reliance on exhaustive candidate evaluations. While it improves planning efficiency, its performance remains limited by the availability of training data and a lack of real-world validation. Similarly, Dai and Zhang [19] take a hybrid approach by integrating machine learning with heuristic optimization, using a dataset of over $0.76$ million measurements to predict signal strength without explicit propagation modeling.
However, existing AI/ML-based methods require extensive labelled datasets and incur substantial data collection and training costs. More importantly, they tend to overlook the sim-to-real discrepancy introduced by inaccurate or incomplete modelling of physical environments, which hinders generalization in real-world deployments.
@@ Acknowledgment
Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the view of the Federal Motor Carrier Safety Administration and/or the U.S. Department of Transportation.

### v3

@@ V-B Automatic Network Planning
The network performance, as defined in Eq. (7), for all compared network planning methods is illustrated in Fig. 7. As we can see, exhaustive search yields the highest overall performance, while random sampling performs significantly worse than both AutoPlan and ES. Notably, AutoPlan achieves performance nearly identical to that of ES, demonstrating its effectiveness in accurate and efficient BS deployment. In terms of computation time, ES requires $602$ minutes, approximately $58.33$ times longer than AutoPlan, highlighting the latter’s practical efficiency.
@@
Moreover, we evaluate the sensitivity of coverage to transmit power in Fig. 7, which plots the coverage achieved by AutoPlan as a function of the number of newly deployed BSs under different transmit power levels. As expected, coverage improves with increased power; however, the gain becomes marginal beyond $30\mathrm{dBm}$. For example, when five new BSs are installed, increasing the power from $30\mathrm{dBm}$ to $43\mathrm{dBm}$ only results in an improvement of $2\%$. Considering the trade-off between performance and power consumption, we adopt $43\mathrm{dBm}$ as the default setting in subsequent experiments. This value yields the highest observed coverage and corresponds to typical configurations of micro-base stations.
@@ VI Related Work
In contrast, AI/ML-based methods aim to enhance scalability and adaptability by replacing manual engineering and handcrafted models with data-driven learning. AutoBS [17] employs reinforcement learning to select BS locations, enabling fast inference during deployment. However, it requires large-scale pretraining on labled environment performance pairs, which is both computationally expensive and time-consuming to collect. OSSN [18], a Transformer-based framework, unifies radio map estimation and site selection to reduce reliance on exhaustive candidate evaluations. While it improves planning efficiency, its performance remains limited by the availability of training data and a lack of real-world validation. Similarly, Dai and Zhang [19] take a hybrid approach by integrating machine learning with heuristic optimization, using a dataset of over $0.76$ million measurements to predict signal strength without explicit propagation modeling.
However, existing AI/ML-based methods require extensive labeled datasets and incur substantial data collection and training costs. More importantly, they tend to overlook the sim-to-real discrepancy introduced by inaccurate or incomplete modeling of physical environments, which hinders generalization in real-world deployments.
@@ Acknowledgment
This work is partially supported by the US National Science Foundation under Grant No. 2321699, and No. 2426481.
This material is based upon work supported by the Federal Motor Carrier Safety Administration under grant number: FM-MHP-0888. Any opinions, findings, and conclusions or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the view of the Federal Motor Carrier Safety Administration and/or the U.S. Department of Transportation.